OpenStudy (anonymous):

|4 + x| - 2x = 7 - 5x

OpenStudy (anonymous):

|4+x| means you can have +(4+x) = +4+x or -(4+x) = -4-x starting from +4+x +4+x-2x=7-5x +x-2x+5x=7-4 find your x with -4-x -4-x-2x=7-5x -x-2x+5x=7+4 find your x it will give you two answers